40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 10 19:09:48 2009, Jenne, Update, PSL, PMC LO Calibration RFSliderAdjustCalib.pngRFSliderAdjustCalibWithOsc.png
    Reply  Sat Apr 11 00:45:41 2009, Yoichi, Update, PSL, PMC LO Calibration 
       Reply  Sun Apr 12 19:27:20 2009, rana, Update, PSL, PMC LO Calibration 
       Reply  Mon Apr 13 08:59:57 2009, steve, Update, PSL, mixers on order 
Message ID: 1471     Entry time: Fri Apr 10 19:09:48 2009     Reply to this: 1473
Author: Jenne 
Type: Update 
Category: PSL 
Subject: PMC LO Calibration 
I measured the RF LO output level from the PMC's LO board which goes directly into the LO input on the PMC Servo board. This goes hand-in-hand with Rana's thoughts
that we might be giving the PMC mixer a too-low LO value, and we might need to switch out the mixer. Steve ordered some new mixers today to try out.

The RF Output Adjust slider (on the C1:PSL_PMC_PS screen) goes from 0-10V; The nominal value (or at least the value I found it at today) is 2.014V.

To measure the RF level: I unlocked the Mode Cleaner and turned off the ISS servo per Yoichi's suggestion. I then unplugged the input to the PMC servo board's LO input,
and put that cable into a 300MHz 'scope, with 12dB attenuation. The 'scope was AC coupled, with the input set to 50Ohms.

I then changed the RF Output Adjust slider in increments of 0.5, and measured the peak-to-peak values on the scope. In the table and on the plots, I've taken into account
the 12dB attenuation. i.e I actually measured 964mV, so 964mV*10^.6 = 3838mV.


RF Output AdjustOutput measured on scopeOscillator Output Monitor
[V]
[Vpp]
[no units given on MEDM screen]
All \pm 0.0159 all of this column is NEGATIVE
0.00003.8380.007
0.50003.8540.007
1.00003.8380.006
1.50003.8380.007
2.00003.8380.006
2.50003.8380.007
3.00003.8380.007
3.50003.8380.007
4.00003.8380.007
4.50003.8220.007
5.00003.8220.012
5.50003.7900.076
6.00003.7580.257
6.50003.6940.555
7.00003.6150.931
7.50003.5351.277
8.00003.4561.532
8.50003.3921.709
9.00003.3441.829
9.50003.3121.908
10.00003.2961.966


I think it's kind of funky that it's so flat for ~half the slider. Also, the third column includes the Oscillator Output Monitor value from the MEDM screen at various RF Adjust slider values. All of these should be negative (i.e. -0.007), but the TABLE function doesn't like "-" signs. I don't know if this information is degenerate with the 'scope measurements, or if it's an indicator of what (might be) wrong.

After finishing, I plugged the cable back into the PMC servo board as it was, turned back on the ISS and relocked the PMC and the MC.
Attachment 1: RFSliderAdjustCalib.png  8 kB  Uploaded Fri Apr 10 20:42:44 2009  | Show | Show all
Attachment 2: RFSliderAdjustCalibWithOsc.png  12 kB  Uploaded Fri Apr 10 20:42:58 2009  | Show | Show all
ELOG V3.1.3-